# SpinMeRound: Consistent multiview Identity Generation Using Diffusion Models

### v1

@@ 3.2 Training Details
Figure 3: Samples generated by SpinMeRound on “in-the-wild” images: Given only the input images (small image in the center), our method produces high-fidelity images from novel angles, along with corresponding shape normals $\mathcal{N}$.
@@ 3.3 Novel view sampling
SpinMeRound synthesizes novel views for an input subject $\mathbf{I}$ given a number of views. In this section, we introduce a robust sampling strategy designed to produce a large number of consistent views that comprehensively cover a full head, given only a single input image. Achieving consistency across viewpoints requires a carefully structured camera pose selection order. Therefore, we employ a three-step sampling process: a) aligning input views and extracting the corresponding shape normals, b) generating anchor images that provide complete coverage of the $360^{\circ}$ human head and c) synthesizing intermediate views by leveraging both the input views and the closest anchor images.
@@
SpinMeRound can generate any arbitrary viewpoint, given the input aligned facial image $\hat{\mathbf{I}}$ and the corresponding shape normals $\mathcal{N}$. However, since it was trained to generate only a limited number of views per sampling process, a carefully designed sampling strategy is essential to produce a wide range of output views. Since the target subject is centered in the scene, we first generate $M=7$ anchor images $\mathbf{A}_{i}$ and corresponding anchor shape normals $\mathbf{A}_{\mathcal{N}_{i}},~{}i\in\{1,...,M\}$, covering a $360^{\circ}$ angle range of the subject (±$\pm 45^{\circ},\pm 90^{\circ},\pm 135^{\circ},180^{\circ}$), as proposed in  [20]. Using these anchors, any number of intermediate views can then be synthesized by conditioning on image triplets $\{(\mathbf{I},\mathcal{N}),(\mathbf{A}_{k},\mathbf{A}_{\mathcal{N}_{k}}),(\mathbf{A}_{l},\mathbf{A}_{\mathcal{N}_{l}})\}$, where $k,l$ where represent the closest anchor images. This approach ensures that each intermediate view remains consistent with both the input aligned image $\hat{\mathbf{I}}$, and the closest already generated views. For both sampling processes, we use the EDM sampler facilitated by the EDM discretization steps [27], while it runs for 50 steps with a guidance scale set to 3. This method enables the generation of 48, 88, or more novel views for a single input subject $\mathbf{I}$, depending on the chosen angle step, thereby covering the entire scene.
Figure 4: Qualitative comparison between SpinMeRound, Zero123-XL [12], SV3D [57] and DiffPortrait3D [24] at angles { ±$\pm 45^{\circ},+180^{\circ}$}. It is clear that SpinMeRound effectively generates high-quality novel views from “in-the-wild” input images, whilst SV3D produces distorted outputs, Zero123-XL generates unnaturally squared avatars and DiffPortrait3D cannot handle such angles.
@@ 4.1 Training Dataset
Training SpinMeRound requires a large multiview dataset containing a large number of subjects $\mathcal{S}$. For each person $\mathcal{S}_{i}$, it is necessary to acquire a set of images $\mathbf{I}_{k}^{i}$, their corresponding shape normal maps $\mathcal{N}_{k}^{i}$, camera poses $\mathcal{C}_{k}^{i}$ alongside their corresponding identity embedding vector $\mathbf{w}^{i}$, where $k\in\{1,2,...,N_{i}\}$ and $N_{i}$ is the number of the available views for the $i$-th scene. Due to the lack of such a public dataset, we create a large-scale synthetic dataset using the publicly available Panohead [1]. We first sample $\sim$10k subjects and manually remove instances with artifacts in the back of the head, resulting in $\sim$7k distinct identities. We render images from 125 different viewpoints for each person to cover the entire head. Simultaneously, we obtain each subject’s facial shape by extracting their opacity values from the triplane feature maps and then applying the marching cubes [38] algorithm. We render their respective normal maps using the acquired facial shape using Pytorch3d [49]. All the images depicting a frontal head are fed to the ArcFace [13] network to extract their corresponding identity vectors $\mathbf{w}$. All in all, we end up with a synthetic dataset containing  7k individuals, rendered from $N=125$ different angles, the respective shape normals, camera poses, and their identity embedding vector $\mathbf{w}$.
@@ 4.2.1 Novel View Synthesis comparisons
In this section, we present a qualitative comparison between our model and other state-of-the-art multiview diffusion models, SV3D [57] Zero123-XL [12] and DiffPortrait3D [24] focusing on their output under ±$\pm 45^{\circ},+180^{\circ}$ angles. SV3D, a latent video diffusion model, creates $360^{\circ}$ videos from a single input image. Zero123-XL, a multiview diffusion model, synthesizes novel perspectives based on the input image and specified camera poses. DiffPortrait3D [24] generates novel views by conditioning the desired camera on an input image and refining the initial noise through a fine-tuning step to achieve optimal results. Figure 4 demonstrates scenarios where the input “in-the-wild” images (Fig. 4a ) are fed to these 4 different networks and their respective generations under ±$\pm 45^{\circ},+180^{\circ}$ viewpoints. As illustrated, SpinMeRound efficiently generates accurate novel viewpoints of the input subject, while SV3D distorts the input image in its generated views, Zero123-XL produces unnatural, square-like facial avatars and DiffPortrait3D cannot handle those angles at all.
Figure 5: Given the input identities (shown in the small squares), we present the results of novel view synthesis after applying 3D Gaussian splatting [28] to the views generated by our model.
@@
To evaluate our model’s consistency in generating novel views from an input image, we assess its ability to reconstruct an input identity through Gaussian Splatting (3DGS) [28]. More specifically, given an input identity, we apply the sampling strategy outlined in Section 3.3 to generate 48 novel views. These views are then used to reconstruct the identity through Gaussian Splatting. As in  [20], we modify the provided 3DGS code to incoporate the LPIPS [65] loss between the ground-truth and the generated viewpoints. In this way, we can deal with small inconsistencies between nearby viewpoints. We present the novel view synthesis of two input identities (small squares) in Figure 5, which clearly illustrates that our proposed methodology can generate consistent subjects.
@@ 4.4 Unconditional Sampling
SpinMeRound is trained following the Classifier-free guidance training scheme. Thus, our architecture can generate novel multiview identities without any prior conditioning. By setting the input identity embedding equal to the empty string, our model can generate a novel identity along with different viewpoints of that identity. Figure 6 presents some examples of those identities.
@@ 4.5 Quantitative Comparisons
We compare our method’s ability to generate novel views with the current state-of-the-art models: a) Eg3D [9] and Panohead [1] which are NeRF-based approaches generating frontal and full-head portraits respectively, b) Zero123 [62], Zero123-XL [12] and DiffPortrai3D [24], which are multiview diffusion models and c) SV3D [57], a multiview video diffusion model. For SV3D, we selected the SV3D-p variant, as it supports flexible viewing angles, whereas SV3D-u operates only with fixed viewpoints. We validate our approach using the NeRSemble dataset [29], which includes 222 unique identities recorded from 16 angles. We randomly select a timestamp from one of their video sequences for each individual and extract a random subset of views. All selected frames are then centered, and one of these views is used as input for the tested methods. We evaluate their reconstruction performance using L2 distance, LPIPS [65], SSIM, and an Identity Similarity score. To calculate the Identity Similarity score, we feed both the ground-truth and the reconstructed images into the ArcFace [13] face recognition network and measure the cosine similarity between their final feature vectors.
Table 1 summarizes the reconstruction performance of each network. SpinMeRound, achieves state-of-the-art results across all cases when evaluated on the LPIPS, SSIM, and Identity Similarity metrics. Additionally, it outperforms all diffusion-based approaches in terms of L2 distance. While our method trails slightly behind NeRF-based approaches, this is expected given the inherent advantages of NeRF-based models in novel view synthesis tasks. Notably, methods like Eg3D and Panohead require a time-intensive fitting process, which can occasionally fail. In contrast, our proposed approach relies solely on an efficient sampling process, eliminating the need for fitting.
@@
NeRF based Eg3d [9] 0.025 0.4 0.55 0.31
Panohead [1] 0.012 0.32 0.65 0.27
Diffusion based Zero123 [36] 0.195 0.515 0.55 0.169
Zero123-XL [12] 0.198 0.51 0.563 0.118
SV3D [57] 0.087 0.41 0.660 0.36
DiffPortrait3d [24] 0.1 0.5 0.35 0.55
SpinMeRound (Ours) 0.033 0.3 0.73 0.61
@@ 5.1 Component analysis
Use of input Image
Starting from a pre-trained Arc2Face model using the EDM framework, we trained a multiview model to generate novel viewpoints given only an input identity embedding. Notably, SpinMeRound has also been trained to be able to generate the input identities without the conditioning view. In this case, our method generates subjects close to the input face, as illustrated in Figure 8.
@@
Identity embedding mechanism
@@
Aiming to showcase the importance of each component, we conducted a multiview reconstruction experiment. More specifically, we sample 100 distinct identities using Panohead [1]. For each subject, we sample 10 different views along with an input frontal view. Given the frontal view as input to each separate model, we reconstruct the remaining views. We measure the performance of each model by calculating their discrepancy using L2 distance, LPIPS [65], SSIM and ID Similarity. As presented in Table 2 and Figure 7, our proposed methodology performs better on all reconstruction metrics while achieving the largest identity similarity score.
@@ 5.2 Importance of ID features
6 Limitations and Future Work
Although SpinMeRound showcases high-fidelity results, it has some limitations. More specifically, our model inherits structural and qualitative limitations from Panohead [1], as our synthetic dataset is derived from it. This means that there are some inconsistencies in the generated eyes, hair and noses. Moreover, the reliance on the alignment step introduces additional potential failure cases, as such errors can propagate through the pipeline.
All in all, the fact that we do not use any capture data limits our model’s capabilities. Hence, this is a direction that we plan to explore in future work. Captured data sets such as FaceScape [68] and NeRSemble [29] could be used to further improve our results. Finally, integrating video diffusion models [5] can be another direction for our future work, to improve the consistency of the generated viewpoints.
@@ 7 Conclusion
In this paper, we presented SpinMeRound, a multiview diffusion model, which generates all-around head portraits of an input subject given a number of input views. Additionally, we introduced a sampling strategy for generating consistent intermediate views, given an unconstrained input “in-the-wild” facial image. Being trained solely on multiview synthetic data, we showcase our method’s abilities by beating the current state-of-the-art multiview diffusion models in novel view synthesis experiments,
Acknowledgments:
@@
S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1) and EPSRC Project GNOMON (EP/X011364/1). B. Kainz received support from the ERC, project MIA-NORMAL 101083647, DFG 512819079, and by the State of Bavaria (HTA). HPC resources were provided by the Erlangen National High Performance Computing Center (NHR@FAU) of the Friedrich-Alexander-Universität Erlangen-Nürnberg (FAU) under the NHR project b180dc. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) - 440719683.
Supplementary Material
Input Panohead [1] Ours
Figure 10: We compare the generated backhead between SpinMeRound(Ours) and Panohead [1].
8 Training Details
SpinMeRound begins training using the publicly available Arc2Face model  [44]. The Arc2Face model is built upon Stable Diffusion 1.5 [51], meaning that it incorporates the following preconditioning functions, according to the EDM framework [27]:
$c_{skip}^{SD1.5}(\sigma)=1,$ $c_{out}^{SD1.5}(\sigma)=-\sigma,$
$c_{in}^{SD1.5}=\frac{1}{\sqrt{\sigma^{2}+1}},$ $c_{noise}^{SD1.5}(\sigma)=\operatorname*{arg\,max}_{j\in[1000]}(\sigma-\sigma_{j})$
As proposed in [27], we modify the aforementioned pre-conditioning by:
$c_{skip}(\sigma)=(\sigma^{2}+1),$ $c_{out}(\sigma)=\frac{-\sigma}{\sqrt{\sigma^{2}+1}},$
$c_{in}=\frac{1}{\sqrt{\sigma^{2}+1}},$ $c_{noise}(\sigma)=0.25\log\sigma,$
Furthermore, we use the proposed noise distribution and weighting functions $log\sigma\sim\mathcal{N}(P_{mean},P^{2}_{std})$ and $\lambda(\sigma)=(1+\sigma^{2})\sigma^{-2}$, with $P_{mean}=0.7$ and $P_{std}=1.6$. We finetune the pre-trained Arc2Face model for 31k iterations, using the training dataset provided by the Arc2Face authors.
8.1 Shape Normals Retrieving
Algorithm 1 Shape Normals sampling using Guidance
1:The aligned facial “in-the-wild” image $\bar{\mathbf{I}}$, the gradient scale $\alpha$, the binary visibility mask $m$, the conditioning mechanism $\mathcal{C}$, and encoder $\mathcal{E}$.
2:$\mathbf{c}\leftarrow\mathcal{C}(\bar{\mathbf{I}})$, $\mathbf{z}_{gt}\leftarrow\{\mathcal{E}(\bar{\mathbf{I}})|\mathbf{0}\}$
3:$\mathbf{z}_{0}\sim\mathcal{N}(\mathbf{0},t_{0}^{2}\mathbf{I})$
4:for all  i from 0 to N-1 do
5:       $\mathbf{\epsilon}_{i}\sim\mathcal{N}(\mathbf{0},S_{noise}^{2}\mathbf{I})$
6:       $\gamma_{i}=\left\{\begin{array}[]{cc}\min(\frac{S_{churn}}{N},\sqrt{2}-1)&\text{if }t_{i}\in[S_{tmin},S_{tmax}]\\ 0&\text{otherwise}\\ \end{array}\right\}$
7:       $\hat{t}_{i}\leftarrow t_{i}+\gamma_{i}t_{i}$
8:       $\hat{\mathbf{x}}_{i}\leftarrow\mathbf{x}_{i}+\sqrt{\hat{t}_{i}^{2}-t_{i}^{2}}\epsilon$
9:       $\mathcal{L}\leftarrow||(\mathbf{z}_{gt}-D_{\theta}(\hat{\mathbf{x}}_{i};\hat{t}_{i},\mathbf{c}))\odot m||_{2}^{2}$
10:       $\mathbf{d}_{i}\leftarrow(\hat{\mathbf{x}}_{i}-D_{\theta}(\hat{\mathbf{x}}_{i};\hat{t}_{i},\mathbf{c})-\alpha\frac{\partial\mathcal{L}}{\partial\hat{\mathbf{x}}_{i}})/\hat{t}_{i}$
11:       $\mathbf{x}_{i+1}\leftarrow\hat{\mathbf{x}}_{i}+(t_{i+1}-\hat{t}_{i})\mathbf{d}_{i}$
12:end for
13:return $\mathbf{z}_{N}$
Input Eg3D [9] Ours
Figure 11: We compare SpinMeRound(Ours) and Eg3D [9] under $+90^{\circ}$ angle.
$-23^{\circ}$ $-16^{\circ}$ $-9^{\circ}$ Input $+9^{\circ}$ $+16^{\circ}$ $+23^{\circ}$
Figure 12: We showcase samples under {±$\{\pm 9^{\circ},\pm 16^{\circ},\pm 23^{\circ}\}$ elevation and azimuth angles.
As mentioned in Section 3.3, given an input “in-the-wild” facial image, we first extract the respective shape normals $\mathcal{N}$. Our proposed sampling methodology is presented in Algorithm  1 and is inspired from Relightify [43]. Given an aligned “in-the-wild” image, we follow the sampling algorithm presented in Algorithm  1, where $\odot$ denotes the Hadamard product $\bar{\mathbf{I}}$. We guide the sampling process to generate the respective shape normals, based on the distribution of the training data. In detail, we firstly extract the conditioning label, as described in Section 3.1 and the latent feature maps of the image $\bar{\mathbf{I}}$, which gets padded, following by sampling the input gaussian noise. For each sampling step, we estimate the $\hat{x}_{i}$ as presented in steps 4, 5, 6 and 7. Then, we compute the guidance loss by calculating the masked L2-distance between the ground-truth latent vector $\mathbf{z}_{gt}$ and the estimated $D_{\theta}(\hat{\mathbf{x}}_{i};\hat{t}_{i},\mathbf{c})$. We calculate the Euler step from $\hat{t}_{i}$ to $t_{i+1}$ by applying the formula in line 9. During sampling we set the guidance scale equal with $10^{4}$ and we run for $t=50$ steps. We set $S_{churn}=0,S_{tmin}=0.05,S_{tmax}=50,S_{noise}=1.003$ and we use the DDPM [51] discretization steps.
9 Qualitative comparison with Panohead and Eg3D
Panohead [1] is a NeRF-based method capable of generating $360^{\circ}$ views. Given an input facial image, it requires a fitting process to produce novel views, often necessitating additional pivotal tuning. In contrast, SpinMeRound eliminates the need for any fitting or fine-tuning steps. Additionally, as presented in Fig. 10, Panohead frequently introduces artifacts on the back of the head, a limitation our method overcomes. On the other hand, EG3D [9] is another NeRF-based method having similar drawbacks as Panohead. Moreover, it only focuses on generating near-frontal views contrary to our full-head approach as shown in Fig 11.
10 Identity sampling
As mentioned in Section 5.1 and presented in Figure 8 of the main paper, our method can generate multiview human identities, given only the input embedding. Although, this work does not focus on multiview identity sampling, we explore our method’s capabilities in this section.
As SpinMeRound has been trained using the classifier-free guidance (CFG)[25] whilst getting 0, 1 or 3 conditioning input images, it can be used to conditionally generate novel images depicting a similar identity as the input one. By setting the guidance scale equal with 3.5, we run the EDM sampler [27] for 50 sampling steps. We set $S_{churn}=0,S_{tmin}=0.05,S_{tmax}=50,S_{noise}=1.003$ and we use the EDM [27] discretization steps, with maximum sigma equal to 700. The sampling process takes about 10sec while it runs on an NVIDIA A100-PCIE. We present samples generated from our model in Figure 13.
11 More samples
We provide additional results in Figures  12,  14 and 15. In Figure 12, we showcase samples generated while using SpinMeRound, under {±$\{\pm 9^{\circ},\pm 16^{\circ},\pm 23^{\circ}\}$ elevation and azimuth angles. Additionally, samples produced from our model are presented in Figures  14 and 15, given the input images on the left. As illustrated, our proposed methodology can be applied to a wide variety of images, including diverse identities, input angles and image styles.
Figure 13: Samples generated using SpinMeRound using only the input identity vector.
Input Images Generated samples
Figure 14: Samples generated with our method, using the images on the left as input (1/2).
Input Images Generated samples
Figure 15: Samples generated with our method, using the images on the left as input (2/2).
Input Images Generated samples

### v2

@@ 3.3 Novel view sampling
SpinMeRound synthesizes novel views for an input subject $\mathbf{I}$, given a number of views. In this section, we introduce a robust sampling strategy designed to produce a large number of consistent views that comprehensively cover a full head, given only a single input image. Achieving consistency across viewpoints requires a carefully structured camera pose selection order. Therefore, we employ a three-step sampling process: a) aligning input views and extracting the corresponding shape normals, b) generating anchor images that provide complete coverage of the $360^{\circ}$ human head and c) synthesizing intermediate views by leveraging both the input views and the closest anchor images.
@@
SpinMeRound can generate any arbitrary viewpoint, given the input aligned facial image $\hat{\mathbf{I}}$ and the corresponding shape normals $\mathcal{N}$. However, since it was trained to generate only a limited number of views per sampling process, a carefully designed sampling strategy is essential to produce a wide range of output views. Since the target subject is centered in the scene, we first generate $M=7$ anchor images $\mathbf{A}_{i}$ and corresponding anchor shape normals $\mathbf{A}_{\mathcal{N}_{i}},~i\in\{1,...,M\}$, covering a $360^{\circ}$ angle range of the subject ($\pm 45^{\circ},\pm 90^{\circ},\pm 135^{\circ},180^{\circ}$), as proposed in  [21]. Using these anchors, any number of intermediate views can then be synthesized by conditioning on image triplets $\{(\mathbf{I},\mathcal{N}),(\mathbf{A}_{k},\mathbf{A}_{\mathcal{N}_{k}}),(\mathbf{A}_{l},\mathbf{A}_{\mathcal{N}_{l}})\}$, where $k,l$ where represent the closest anchor images. This approach ensures that each intermediate view remains consistent with both the input aligned image $\hat{\mathbf{I}}$, and the closest already generated views. For both sampling processes, we use the EDM sampler facilitated by the EDM discretization steps [31], while it runs for 50 steps with a guidance scale set to 3. This method enables the generation of 48, 88, or more novel views for a single input subject $\mathbf{I}$, depending on the chosen angle step, thus covering the entire scene.
Figure 4: Qualitative comparison between SpinMeRound, Zero123-XL [13], SV3D [65] and DiffPortrait3D [26] at angles $\pm 45^{\circ},+180^{\circ}$}. It is clear that SpinMeRound effectively generates high-quality novel views from “in-the-wild” input images, whilst SV3D produces distorted outputs, Zero123-XL generates unnaturally squared avatars and DiffPortrait3D cannot handle such angles.
@@ 4.1 Training Dataset
Training such a model requires a large multi-view dataset containing a great number of subjects $\mathcal{S}$. For each person $\mathcal{S}_{i}$, it is necessary to acquire a set of images $\mathbf{I}_{k}^{i}$, their corresponding shape normal maps $\mathcal{N}_{k}^{i}$, camera poses $\mathcal{C}_{k}^{i}$ alongside with their corresponding identity embedding vector $\mathbf{w}^{i}$, where $k\in\{1,2,...,N_{i}\}$ and $N_{i}$ is the number of the available views for the $i$-th scene. Due to the lack of such a dataset, we create a large-scale synthetic dataset using the publicly available Panohead [1]. We first sample $\sim$10k subjects and manually remove instances with artifacts in the back of the head, resulting in $\sim$7k distinct identities. We render images from 125 different viewpoints for each person to cover the entire head. Simultaneously, we obtain each subject’s facial shape by extracting their opacity values from the triplane feature maps and then applying the marching cubes [42] algorithm. We render their respective normal maps using the acquired facial shape using Pytorch3d [56]. All the images depicting a frontal head are fed to the Arc-Face [14] network to extract their corresponding identity vectors $\mathbf{w}$. All in all, we end up with a synthetic dataset containing about 7k individuals, rendered from $N=125$ different angles, the respective shape normals, camera poses, and their identity embedding vector $\mathbf{w}$.
@@ 4.2.1 Novel View Synthesis comparisons
In this section, we present a qualitative comparison between our model and other state-of-the-art multi-view diffusion models, SV3D [65] Zero123-XL [13] and DiffPortrait3D [26], focusing on their output under $\pm 45^{\circ},+180^{\circ}$ angles. SV3D, a latent video diffusion model, creates $360^{\circ}$ videos from a single input image. Zero123-XL, a multi-view diffusion model, synthesizes novel perspectives based on the input image and specified camera poses. DiffPortrait3D [26] generates novel views by conditioning the desired camera on an input image and refining the initial noise through a fine-tuning step to achieve optimal results. Figure 4 demonstrates scenarios where the input “in-the-wild” images (Fig. 4a ) are fed to these 4 different networks and their respective generations under $\pm 45^{\circ},+180^{\circ}$ viewpoints. As illustrated, SpinMeRound efficiently generates accurate novel viewpoints of the input subject, while SV3D distorts the input image in its generated views, Zero123-XL produces unnatural, square-like facial avatars and DiffPortrait3D cannot handle those angles at all.
Figure 5: Given the input identities (shown in the small boxes), we present the results of novel view synthesis after applying 3D Gaussian splatting [32] to the views generated by our model.
@@
To evaluate our model’s consistency in generating novel views from an input image, we assess its ability to reconstruct an input identity through Gaussian Splatting (3DGS) [32]. More specifically, given an input identity, we apply the sampling strategy outlined in Section 3.3 to generate 48 novel views. These views are then used to reconstruct the identity through Gaussian Splatting. As in  [21], we modify the provided 3DGS code to incorporate the LPIPS [72] loss between the ground-truth and the generated viewpoints. In this way, we can deal with small inconsistencies between nearby viewpoints. We present the novel view synthesis of two input identities (small squares) in Figure 5, which clearly illustrates that our proposed methodology can generate consistent subjects.
@@ 4.4 Unconditional Sampling
SpinMeRound is trained following the Classifier-Free Guidance training scheme. Thus, our architecture can generate novel multi-view identities without any prior conditioning. By setting the input identity embedding equal to the empty string, our model can generate a novel identity along with different viewpoints of that identity. Figure 6 presents some examples of those identities.
@@ 4.5 Quantitative Comparisons
We compare our method’s ability to generate novel views with the current state-of-the-art models: a) Eg3D [9] and Panohead [1] which are NeRF-based approaches generating frontal and full-head portraits respectively, b) Zero123 [11], Zero123-XL [13] and DiffPortrai3D [26], which are multi-view diffusion models and c) SV3D [65], a multi-view video diffusion model. For SV3D, we selected the SV3D-p variant, as it supports flexible viewing angles, whereas SV3D-u operates only with fixed viewpoints. We validate our approach using the NeRSemble dataset [33], which includes 222 unique identities recorded from 16 angles. We randomly select a timestamp from one of their video sequences for each individual and extract a random subset of views. All selected frames are then centered, and one of these views is used as input for the tested methods. We evaluate their reconstruction performance using L2 distance, LPIPS [72], SSIM, and an Identity Similarity score. To calculate the Identity Similarity score, we feed both the ground-truth and the reconstructed images into the ArcFace [14] and VGGFace [50] networks and measure the cosine similarity between their final feature vectors.
Table 1 summarizes the reconstruction performance of each network. SpinMeRound, achieves state-of-the-art results across all cases when evaluated on the LPIPS, SSIM, and Identity Similarity metrics. Additionally, it outperforms all diffusion-based approaches in terms of L2 distance. Although our method trails slightly behind NeRF-based approaches, this is expected given the inherent advantages of NeRF-based models in novel view synthesis tasks. Notably, methods like Eg3D and Panohead require a time-intensive fitting process, which can occasionally fail. In contrast, our proposed approach relies solely on an efficient sampling process, eliminating the need for fitting.
@@
NeRF based Eg3d [9] 0.025 0.4 0.55 0.31 0.89
Panohead [1] 0.012 0.32 0.65 0.27 0.88
Diffusion based Zero123 [40] 0.195 0.515 0.55 0.169 0.44
Zero123-XL [13] 0.198 0.51 0.563 0.118 0.442
SV3D [65] 0.087 0.41 0.660 0.36 0.881
DiffPortrait3d [26] 0.1 0.5 0.35 0.55 0.887
SpinMeRound (Ours) 0.033 0.3 0.73 0.61 0.911
@@ 5.1 Component analysis
Use of input image:
Starting from a pre-trained Arc2Face model using the EDM framework, we trained a multi-view model to generate novel viewpoints given only an input identity embedding vector. Notably, SpinMeRound has also been trained to be able to generate input identities without providing the conditioning view. In this case, it generates subjects close to the input face, as illustrated in Figure 8.
@@
Identity embedding mechanism:
@@
Aiming to showcase the importance of each component, we conducted a multi-view reconstruction experiment. More specifically, we sample 100 distinct identities using Panohead [1]. For each subject, we sample 10 different views along with an input frontal view. Given the frontal view as input to each separate model, we reconstruct the remaining views. We measure the performance of each model by calculating their discrepancy using L2 distance, LPIPS [72], SSIM and ID Similarity. As presented in Table 2 and Figure 7, our proposed methodology performs better regarding all reconstruction metrics while achieving the highest identity similarity score.
@@ 6 Conclusion
In this paper, we presented SpinMeRound, a multi-view latent diffusion model, which generates all-around head portraits of an input subject given a number of input views alongside an input identity embedding. Additionally, we introduced a sampling strategy for generating consistent intermediate views, given an unconstrained input “in-the-wild” facial image. Being trained sorely on multi-view synthetic data, we showcase our method’s abilities by beating the current state-of-the-art multi-view diffusion models in novel view synthesis experiments.
@@
Acknowledgments: S. Zafeiriou and part of the research was funded by the EPSRC Fellowship DEFORM (EP/S010203/1) and EPSRC Project GNOMON (EP/X011364/1). The authors gratefully acknowledge the scientific support and HPC resources provided by the Erlangen National High Performance Computing Center (NHR@FAU) of FAU under the NHR projects b143dc and b180dc. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the German Research Foundation (DFG) – 440719683. Additional support was also received by the ERC - project MIA-NORMAL 101083647, DFG 513220538, 512819079, and by the state of Bavaria (HTA).